what is the formula for solving equation -50t+10=35?
Answer by AnlytcPhil(1806)

                         -50t + 10 = 35

There are two terms on the left and one term on the right.

Use addition or subtraction of the same quantity on both sides to:

1. Get rid of all terms on the left that DO NOT contain the variable

2. Get rid of all terms on the right that DO contain the variable

                         -50t + 10 = 35

The first term -50x contains the variable and is on the left, so we do
not want to get rid of it.  The next term on the left is +10.  It is on
the left side and DOES NOT contain the variable, so we need to get rid of
the +10 on the left, by adding it's opposite -10 to both sides.

                         -50t + 10 =  35
                               -10   -10
                      ————————————————————
                         -50t      =  25

3. When there is only one term with a avariable on the left, and only
one term on the right without a variable, then divide both sides by the
coefficient ot the variable.

So we divide both sides by -50

                         -50t       25
                        ——————— = —————    
                         -50       -50                 

The left sides reduces to just t, and the right side reduces to -1/2

                             t  =  -1/2

Edwin
AnlytcPhil@aol.com